# Fusing Radiomic Features with Deep Representations for Gestational Age Estimation in Fetal Ultrasound Images

### v1

@@ 2.2 Deep Representation Learning
When working with a grayscale fetal ultrasound image $X_{\text{US}}$, the initial step involves converting $X_{\text{US}}$ into a 3-channel image denoted as $X_{\text{US}}\in\mathbb{R}^{3\times n_{\text{h}}\times n_{\text{w}}}$, where $n_{\text{h}}$ and $n_{\text{w}}$ represent the height and width of the image, respectively. This conversion is essential to facilitate the fine-tuning of pre-trained DL models designed to process 3-channel input images. The input $X_{\text{US}}$ is processed by the model $f_{\theta}$, producing a deep representation vector that is subsequently fed to the $\operatorname{ReLU}$ activation and linearly combined to output $\bm{x}_{\text{DL}}\in\mathbb{R}^{512}$. Subsequently, $\bm{x}_{\text{DL}}$ is fused with radiomic features $\bar{\bm{x}}_{\text{RAD}}$ within the cross-attention module.
@@ 2.3 Radiomic Features
Our dataset comprises fetal head ultrasound images and corresponding region of interest (ROI) masks, denoted as $X_{\text{US}}$ and $X_{\text{ROI}}$, respectively. The ROI encompasses the full extent of the fetal head, as annotated by sonographers. The input $X_{\text{ROI}}$ has dimensions of $1\times n_{\text{h}}\times n_{\text{w}}$, where $n_{\text{h}}$ and $n_{\text{w}}$ correspond to the height and width of ROI image. The process of extracting radiomic features utilizes each pair of $X_{\text{US}}$ and $X_{\text{ROI}}$ as input. In this study, a total of 95 radiomic features $\bm{x}_{\text{RAD}}\in\mathbb{R}^{95}$, which capture surface-level details of the fetal head region, are extracted with the publicly accessible Python package, pyradiomics [^1] https://pyradiomics.re adthedocs.io/en/latest/. These extracted features are categorized into shape, statistical (first-order), and texture features [7, 2, 9].
@@ 2.5 Model Training and Loss Functions
We employ a fine-tuning strategy for our framework, initializing the DL models with pre-trained weights from ImageNet [5]. Only the trainable components in our proposed framework are fine-tuned, see Fig. 1. We optimize the likelihood of the $\hat{y}$ regression output and therefore use the least squares for the overall training loss: $\mathcal{L}=\sum_{i=1}^{M}\left(y^{\langle i\rangle}-\hat{y}^{\langle i\rangle}\right)^{2}$, where $\hat{y}^{\langle i\rangle}$ is the model predicted GA for the $i^{\text{th}}$ training input and $y^{\langle i\rangle}$ is the computed GA as described in Section 3.1. $M$ is the size of the training set. The model with the best performance is saved for testing based on their predictive accuracy during the model training phase.
@@ 3.1 Datasets
We utilize two public datasets in this study: the Spanish trans-thalamic dataset (ES-TT) originates two centers in Spain [4]; the HC18 dataset is sourced from a database in Netherlands [12]. All scans were acquired from healthy singleton pregnancies. HC18 data was collected using two ultrasound machines from a single center, while ESTT data was obtained from six machines across two different centers. The variability of the fetal head across various pregnancy trimesters poses a challenge in estimating GA. We combine two datasets into one and use a 70:30 ratio to split the data into training and testing sets. A subset of 100 images from the test set is randomly selected for validation purposes.
@@ 3.2 Implementation Details
The input image $X_{\text{US}}$ and mask $X_{\text{ROI}}$ are resized to the same resolution, $256\times 256$. To decrease computational demands, $X_{\text{US}}$ is normalized following the same process from [11]. The Adam optimizer is utilized with an initial learning rate of $10^{-5}$ following a weight decay of $10^{-6}$. The batch size is fixed at 8, and the epoch is set to 60. Our framework is implemented using Pytorch library and trained on a single RTX 4090 GPU with 24 GB memory.
@@ 3.3 Results
To assess the accuracy of predicting GA, we use Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) as the evaluation metrics. We find that the correlation coefficient between MAE and RMSE is larger than 0.998, therefore, we only report the MAE in this paper. We evaluate our method by comparing it with Baseline utilizing only images and Concatenation utilizing concatenated image and radiomic features.

### v2

@@ 2.2 Deep Representation Learning
The input $X_{\text{US}}\in\mathbb{R}^{3\times n_{\text{h}}\times n_{\text{w}}}$ is processed by the model $f_{\theta}$, producing a deep representation vector that is subsequently fed to the $\operatorname{ReLU}$ activation and linearly combined to output $\bm{x}_{\text{DL}}\in\mathbb{R}^{512}$. Subsequently, $\bm{x}_{\text{DL}}$ is fused with radiomic features $\bar{\bm{x}}_{\text{RAD}}$ within the cross-attention module.
@@ 2.3 Radiomic Features
Our dataset comprises fetal head ultrasound images and corresponding region of interest (ROI) masks, denoted as $X_{\text{US}}$ and $X_{\text{ROI}}$, respectively. The input $X_{\text{ROI}}$ has dimensions of $1\times n_{\text{h}}\times n_{\text{w}}$, where $n_{\text{h}}$ and $n_{\text{w}}$ correspond to the height and width of ROI image. The process of extracting radiomic features utilizes each pair of $X_{\text{US}}$ and $X_{\text{ROI}}$ as input. In this study, a total of 95 radiomic features $\bm{x}_{\text{RAD}}\in\mathbb{R}^{95}$, which capture surface-level details of the fetal head region, are extracted with the publicly accessible Python package, pyradiomics [^1] https://pyradiomics.re adthedocs.io/en/latest/. These extracted features are categorized into shape, statistical (first-order), and texture features [7, 2, 9].
@@ 2.5 Model Training and Loss Functions
We employ a fine-tuning strategy for our framework, initializing the DL models with pre-trained weights from ImageNet [5]. We optimize the likelihood of the $\hat{y}$ regression output and therefore use the least squares for the overall training loss: $\mathcal{L}=\sum_{i=1}^{M}\left(y^{\langle i\rangle}-\hat{y}^{\langle i\rangle}\right)^{2}$, where $\hat{y}^{\langle i\rangle}$ is the model predicted GA for the $i^{\text{th}}$ training input and $y^{\langle i\rangle}$ is the computed GA as described in Section 3.1. $M$ is the size of the training set. The model with the best performance is saved for testing based on their predictive accuracy during the model training phase.
@@ 3.1 Datasets
We utilize two public datasets in this study: the Spanish trans-thalamic dataset (ES-TT) originates two centers in Spain [4]; the HC18 dataset is sourced from a database in Netherlands [12]. All scans were acquired from healthy singleton pregnancies. The variability of the fetal head across various pregnancy trimesters poses a challenge in estimating GA. We use a 70:30 ratio to split the data into training and testing sets. A subset of 100 images from the test set is randomly selected for validation purposes.
@@ 3.2 Implementation Details
The input image $X_{\text{US}}$ and mask $X_{\text{ROI}}$ are resized to the same resolution, $256\times 256$. The Adam optimizer is utilized with an initial learning rate of $10^{-5}$ following a weight decay of $10^{-6}$. The batch size is fixed at 8, and the epoch is set to 60. Our framework is implemented using Pytorch library and trained on a single RTX 4090 GPU with 24 GB memory.
@@ 3.3 Results
To assess the accuracy of predicting GA, we use Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) as the evaluation metrics. We find that the correlation coefficient between MAE and RMSE is larger than 0.998, therefore, we only report the MAE in this paper.